# Offline Dedicated Quantum Attacks on Block Ciphers Based on Two Parallel Permutation-Based Pseudorandom Functions

### v2

@@ 3.2.1 Instantiation of TPP-PRFs
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}x^{l_{1}}\oplus l_{14}k_{1}^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{24}k_{2}^{l}\|0^{t}\oplus u)$
$\oplus l_{33}P_{1}(l_{13}x^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)\oplus C$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{14}k_{1})\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus l_{24}k_{2})$
$\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)\right)\oplus l_{34}P_{2}\left(l_{23}(x^{l_{1}}\|0^{n-p-t})\oplus l_{24}k_{2}^{l}\|0^{t}\oplus u\right)\oplus C$
@@
Next, we demonstrate that several instantiations of TPP-PRFs can construct $p$-XOR-type function, including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
The Xop construction [1] is defined as the bitwise XOR of the outputs from two distinct pseudorandom permutations (PRPs) applied to the same input $x$:
@@
The Xop construction is realized with two Even-Mansour ciphers, $EM_{1}(x)=P_{1}(x\oplus k_{1})\oplus k_{2}$ and $EM_{2}(x)=P_{2}(x\oplus k_{3})\oplus k_{4}$, resulting in the $\mathsf{XopEM}$ function.
@@
We can obtain the $p$-XOR-type function $f(i,x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus P_{1}(x)\oplus P_{2}(x\oplus i)\oplus k_{2}\oplus k_{4}$, where $g_{1}(x)=\mathsf{XopEM}(x)$, a public function $p(i,x)=P_{1}(x)\oplus P_{2}(x\oplus i)$. This function has a period $s=k_{1}$ precisely when $i=k_{3}$.
@@
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})$ $=\bigoplus\limits_{u\in\mathcal{L}}P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
Consider a periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$,
@@
SoEM22 [7]. The Sum of Even-Mansour (SoEM) construction employs two public permutations $P_{1}$, $P_{2}$, and two $n$-bit keys $k_{1}$, $k_{2}$. It represents a specific instance of the XOR construction where $k_{3}=k_{2}$ and $k_{4}=k_{1}$, defined as:
@@
We can construct $p$-XOR-type function of the form $f(i,x)=\mathsf{SoEM22}(x)\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$, where $P_{1}(x)\oplus P_{2}(x\oplus i)$ is a public function, which satisfies $f(i,x)=f(i,x\oplus k_{1})$ when $i=k_{2}$. The new periodic function is written as $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u)$, which has a period $k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=k_{2}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{2}^{l}\|k_{1}^{l_{2}}$.
Consider $P_{1}=P$ and $P_{2}=P^{-1}$, the above $\mathsf{SoEM22}$ construction corresponds to $\mathsf{SUMPIP}$ [12]. In the same way, it can construct a $p$-XOR-type function $f(i,x)$ and generate a new periodic function $F^{\mathcal{L}}$ based on the truncation parameter.
DS-SoEM [2]. This construction is a domain-separated variant of the Xop construction, which is a sum of two Even-Mansour ciphers using a $d$-bit domain-separation constant. For an input $x\in\{0,1\}^{n-d}$, the encryption is defined as:
$\mathsf{DS\text{-}SoEM}(x)=$ $P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{1})\right)\|0^{d}\right)\oplus P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{2})\right)\|1^{d}\right)\oplus k_{1}\oplus k_{2},$ (13)
@@
We define a $p$-XOR-type function $f(i,x)=\mathsf{DS\text{-}SoEM}(x)\oplus P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$, which satisfies $f(i,x)=f\left(i,x\oplus\mathrm{msb}_{n-d}(k_{1})\right)$ when $i=\mathrm{msb}_{n-d}(k_{2})$, where $P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$ is a public function. Then, we define the following function
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus_{u\in\mathcal{L}}\mathsf{DS\text{-}SoEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)\oplus P\left((x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\|0^{d}\right)$ (14)
@@
This function exhibits periodicity, that is, $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus[\mathrm{msb}_{n-d}(k_{1})]^{l_{1}})$ when $i^{l}\|j^{l_{2}}=[\mathrm{msb}_{n-d}(k_{2})]^{l}\|[\mathrm{msb}_{n-d}(k_{1})]^{l_{2}}$.
3.2.2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
Based on the above $p$-XOR-type function constructed by TPP-PRFs, we now propose an offline dedicated quantum attack on them, utilizing the truncated technique in the Q1 (Q2) model, which reduces query complexity to the encryption oracle.
We present a new algorithm, termed the offline dedicated quantum attack on TPP-PRFs-based construction, which integrates the offline Simon’s algorithm with the dedicated quantum attack on $p$-XOR-type function. The main idea is to separate this attack into two phases: the online phase, which prepares the superposition state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ through classical or quantum queries to the encryption oracle $U_{g_{1}}$ (with classical queries requiring $2^{n}$ in the Q1 model by Algorithm 6, and quantum queries requiring parallel access to $2^{t}$ distinct $U_{g_{1}}$ in the Q2 model by Algorithm 7), and the offline computation phase, which performs the remaining quantum computations independently of the encryption oracle (with the test procedure checking whether the function $F^{\mathcal{L}}$ has a period without any new query to the encryption oracle in Algorithm 5). This separation significantly reduces the query complexity, as Grover iterations are executed offline after a limited number of online interactions.
Now, we introduce the offline dedicated quantum attack on block ciphers based on $p$-XOR-type function in the Q1 (Q2) model by Algorithm 4.
Algorithm 4 The Offline Dedicated Attack on block ciphers based on $p$-XOR-Type Function Using Truncated Technique
@@ 3.2.2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
7:Apply unitary operation to obtain:
@@
The overall quantum attack aims to find the high $(\kappa-t)$ bits $i_{0}^{l}$ such that it $F^{\mathcal{L}}(i_{0}^{l},\cdot)$ exhibits a non-trivial period $s^{l}$. Starting with the superposition state $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$, the input is $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$ of $(\kappa-t)+c^{\prime}(n-t)+c^{\prime}m$ qubits. The initial state $|\Psi\rangle$ required for the Grover iteration is obtained by several $H$ transforms. The iterative function is $G=D_{|\Psi\rangle}O_{\text{test}}$ for $\mathcal{O}(2^{(\kappa-t)/2})$ iterations, where $D_{|\Psi\rangle}=(2|\Psi\rangle\langle\Psi|-I)$. The final step is measuring the index $i^{l}$ to obtain $i_{0}^{l}$. The implementation of $U_{P^{\mathcal{L}}}$ and $O_{\text{test}}$ in Algorithm 4 follows an approach similar to that used in Ref. [23]. We briefly illustrate the implementation of $U_{P^{\mathcal{L}}}$. The function
$P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u)$
is the XOR sum of $2^{t}$ public functions $p$, which can be implemented in parallel without a sequential relationship, where $t$ is a tunable truncation parameter.
To get a usable success probability after $r$ iterations, we need $c^{\prime}=n+\kappa-2t+\pi+1$ queries per iteration. Next, we give the complexity and success probability of the offline dedicated quantum attack on block ciphers based on $p$-XOR-type function in the Q1 (Q2) model by Theorem 4.
Theorem 4.
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a $p$-XOR-type function, then there exists the offline dedicated quantum attack on $f$ by Algorithm 4. This attack can recover $i_{0}^{l}\in\{0,1\}^{\kappa-t}$ with a time complexity of $O(2^{(\kappa-t)/2}\cdot(n^{3}+T_{p}))$, requiring $O(2^{n})$ classical queries to the encryption oracle in the Q1 model, and $O(2^{t})$ quantum queries in the Q2 model. The probability of falling below a success rate of
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $c^{\prime}=n+\kappa-2t+\tau+1$ and $T_{p}$ is the time required to compute the public function $p$ once.
For $p$-XOR-type functions and new periodic functions $F^{\mathcal{L}}$ constructed from some instantiations of TPP-PRFs in Section 3.2.1, we can implement offline dedicated attacks on them to recover the period of $F^{\mathcal{L}}$ by Theorem 4.
@@
In the Q1 model, this attack on block ciphers based on TPP-PRFs requires $O(2^{p+t})$ classical queries to the encryption oracle and $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time. Note that each iteration of $p$ can be completed in $O(1)$ time. Let $D$ and $T$ be the number of classical queries and the time of offline quantum computations. A classical attack exhibits a quantum-time/classical-query tradeoff of $T^{2}D=2^{2n-t}$, which is balanced at $T=D=2^{(2n-t)/3}$ for a tunable truncation parameter $t$.
@@
The remaining keys can be recovered as follows. Taking TPP-PRFs as an example, after we obtain $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$, Simon’s algorithm is used to recover $l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ by requiring $O(2^{p+t})$ classical queries, and its offline computation requires $O(n^{3})$ time. Then, we can easily recover the remaining $t$-bit by applying the offline Simon’s algorithm on $f^{\prime}(k^{r},x)=f(k_{2}^{l}\|i^{r},x)$, which has the period $s$ when $i^{r}=k_{0}^{r}$.
@@
In the Q2 model, consider the offline attack proposed in Theorem 4, the construction of $F^{\mathcal{L}}$ differs from that in Section 3.2.1 in that there is unnecessary to partition the (nt)(n-t)-bit subkey $k_{1}^{l}$ into $k_{1}^{l_{1}}$ with $p$-bit and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bit.
@@
Taking the $\mathsf{SoEM22}$ construction as an example, we can construct $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$ with the public function $P_{2}(x\oplus i)$, then obtain a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})=\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l}\|0^{t}\oplus u)\oplus P_{1}(x^{l}\|0^{t}\oplus u)\oplus P_{2}\left((x^{l}\oplus i^{l})\|0^{t}\oplus u\right)$ that $F^{\mathcal{L}}(i^{l},x^{l})$ has an (nt)(n-t)-bit period $s^{l}=k^{l}_{1}$ when $i^{l}=k^{l}_{2}$. Based on this, we can prepare the state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ by requiring $2^{t}$ quantum queries to $\mathsf{SoEM22}$ by Algorithm 7. Then, we apply the offline attack by Algorithm 4, which successfully recovers the period $k^{l}_{1}$ with $O\left(2^{(n-t)/2}\cdot(n^{3}+T_{p})\right)$ offline quantum computation time. Compared to the attack by Shi et al. [23], our method takes the same time complexity but reduces the quantum query complexity from $O\left(2^{(n+t)/2}\cdot(n-t)\right)$ to $O(2^{t})$, demonstrating the advantage of our offline attack in the Q2 model.
4 Conclusion

### v3

@@ III-B1 Instantiation of TPP-PRFs
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}\left((l_{13}x^{l_{1}}\oplus l_{14}k_{1}^{l_{1}})\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u\right)$
$\oplus l_{34}P_{2}\left((l_{23}x^{l_{1}}\oplus l_{24}k_{2}^{l_{1}})\|l_{24}k_{2}^{l_{2}}\|0^{t}\oplus u\right)\oplus C$
$\oplus l_{33}P_{1}\left(l_{13}(x^{l_{1}}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}})\|0^{t}\oplus u\right)$
$\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)$
$\qquad=$ $\bigoplus\limits_{u\in\mathcal{L}}l_{33}P_{1}(l_{13}x^{l_{1}}\|l_{14}k_{1}^{l_{2}}\|0^{t}\oplus u)$
$\oplus l_{34}P_{2}\left((l_{23}x^{l_{1}}\oplus l_{23}l_{13}^{-1}l_{14}k_{1}^{l_{1}}\oplus l_{24}k_{2}^{l_{1}})\|l_{24}k_{2}^{l_{2}}\|0^{t}\oplus u\right)$
$\oplus l_{33}P_{1}\left((l_{13}(x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}})\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}))\|0^{t}\oplus u\right)\oplus C$
$\oplus l_{34}P_{2}(l_{23}(x^{l_{1}}\oplus l_{13}^{-1}l_{14}k_{1}^{l_{1}})\|0^{n-t-p}\oplus l_{23}^{-1}l_{24}k_{2}^{l})\|0^{t}\oplus u)$
@@
Next, we demonstrate that several instantiations of TPP-PRFs can be used to construct $p$-XOR-type functions, including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$.
The Xop construction [bellare1998luby] is defined by the bitwise XOR of the outputs from two distinct pseudorandom permutations (PRPs) applied to the same input $x$:
@@
Realizing the Xop construction with two Even-Mansour ciphers, $EM_{1}(x)=P_{1}(x\oplus k_{1})\oplus k_{2}$ and $EM_{2}(x)=P_{2}(x\oplus k_{3})\oplus k_{4}$, resulting in the $\mathsf{XopEM}$ function.
@@
We can obtain a $p$-XOR-type function as follows: $f(i,x)=P_{1}(x\oplus k_{1})\oplus P_{2}(x\oplus k_{3})\oplus P_{1}(x)\oplus P_{2}(x\oplus i)\oplus k_{2}\oplus k_{4}$, where $g_{1}(x)=\mathsf{XopEM}(x)$, and the public function is $p(i,x)=P_{1}(x)\oplus P_{2}(x\oplus i)$. This function has a period $s=k_{1}$ precisely when $i=k_{3}$.
@@
$P^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus\limits_{u\in\mathcal{L}}P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)$
$\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
The periodic function $F^{\mathcal{L}}:\{0,1\}^{n-t}\times\{0,1\}^{n-t-p}\times\{0,1\}^{p}\rightarrow\{0,1\}^{n}$ is defined as:
@@
SoEM22 [chen2019build]. The Sum of Even-Mansour (SoEM) construction employs two public permutations $P_{1}$, $P_{2}$, and two $n$-bit keys $k_{1}$, $k_{2}$. It is a specific instance of the XOR construction where $k_{3}=k_{2}$ and $k_{4}=k_{1}$, defined as:
@@
We can construct a $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}(x)\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$, where $P_{1}(x)\oplus P_{2}(x\oplus i)$ is a public function. This function satisfies $f(i,x)=f(i,x\oplus k_{1})$ when $i=k_{2}$. The new periodic function $F^{\mathcal{L}}$ is defined as:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$ (13)
$\oplus P_{1}(x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)$
$\oplus P_{2}(((x^{l_{1}}\|0^{n-t-p})\oplus i^{l})\|0^{t}\oplus u).$
This function has a period $k_{1}^{l_{1}}$ when $i^{l}\|j^{l_{2}}=k_{2}^{l}\|k_{1}^{l_{2}}$ or $i^{l}\|j^{l_{2}}=k_{1}^{l}\oplus k_{2}^{l}\|k_{1}^{l_{2}}$.
If $P_{1}=P$ and $P_{2}=P^{-1}$, the above $\mathsf{SoEM22}$ construction corresponds to $\mathsf{SUMPIP}$ [guo2019beyond]. In the same way, a $p$-XOR-type function $f(i,x)$ and a new periodic function $F^{\mathcal{L}}$ can be constructed based on the truncation parameter.
DS-SoEM [bhattacharjee2022cencpp]. This construction is a domain-separated variant of the Xop construction that uses a $d$-bit domain-separation constant and is a sum of two Even-Mansour ciphers. For an input $x\in\{0,1\}^{n-d}$, it is defined as:
$\mathsf{DS\text{-}SoEM}(x)=$ $P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{1})\right)\|0^{d}\right)$ (14)
$\oplus P\left(\left(x\oplus\mathrm{msb}_{n-d}(k_{2})\right)\|1^{d}\right)\oplus k_{1}\oplus k_{2},$
@@
We define a $p$-XOR-type function $f(i,x)=\mathsf{DS\text{-}SoEM}(x)\oplus P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$, where the public function is $P(x\|0^{d})\oplus P\left((x\oplus i)\|1^{d}\right)$. This function satisfies $f(i,x)=f\left(i,x\oplus\mathrm{msb}_{n-d}(k_{1})\right)$ when $i=\mathrm{msb}_{n-d}(k_{2})$. The new function $F^{\mathcal{L}}$ is defined as:
$F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=$ $\bigoplus_{u\in\mathcal{L}}\mathsf{DS\text{-}SoEM}(x^{l_{1}}\|0^{n-t-p}\|0^{t}\oplus u)$ (15)
$\oplus P\left((x^{l_{1}}\|j^{l_{2}}\|0^{t}\oplus u)\|0^{d}\right)$
@@
This function is periodic, with $F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}})=F^{\mathcal{L}}(i^{l}\|j^{l_{2}},x^{l_{1}}\oplus[\mathrm{msb}_{n-d}(k_{1})]^{l_{1}})$ when $i^{l}\|j^{l_{2}}=[\mathrm{msb}_{n-d}(k_{2})]^{l}\|[\mathrm{msb}_{n-d}(k_{1})]^{l_{2}}$.
Algorithm 4 Offline Dedicated Attack on block cipher constructions based on $p$-XOR-Type Function Using Truncated Technique
@@
III-B2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
Based on the above $p$-XOR-type functions constructed by TPP-PRFs, we propose an offline dedicated quantum attack on these constructions, utilizing the truncated technique in the Q1 (Q2) model to reduce query complexity to the encryption oracle.
We present a new algorithm for an offline dedicated quantum attack on constructions based on TPP-PRFs, which integrates the offline Simon’s algorithm with the dedicated quantum attack on block ciphers by constructing $p$-XOR-type functions. Our algorithm is separated into two phases: the online phase prepares the superposition state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ via classical or quantum queries to the encryption oracle $U_{g_{1}}$ (requiring $O(2^{n})$ classical queries in the Q1 model by Algorithm 6, or $O(2^{t}\cdot(n-t))$ quantum queries in the Q2 model by Algorithm 7), and the offline computation phase performs the remaining quantum computations independently of the encryption oracle (using the test procedure in Algorithm 5 to check whether the function $F^{\mathcal{L}}$ has a period without any new queries to the encryption oracle). This separation achieves a reduction in query complexity, as Grover iterations are executed offline after a limited number of online interactions.
The offline dedicated quantum attack on constructions based on $p$-XOR-type functions in the Q1 (Q2) model is presented in Algorithm 4.
The overall quantum attack aims to find the high $(\kappa-t)$ bits $i_{0}^{l}$ such that it $F^{\mathcal{L}}(i_{0}^{l},\cdot)$ exhibits a non-trivial period $s^{l}$. Starting with the superposition state $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle=\otimes^{c^{\prime}}\left(\sum_{x^{l}\in\{0,1\}^{n-t}}|x^{l}\rangle|G_{1}^{\mathcal{L}}(x^{l})\rangle\right)$, the input is $\left|\psi_{G_{1}^{\mathcal{L}}}\right\rangle|0^{\kappa-t}\rangle$ of $(\kappa-t)+c^{\prime}(n-t)+c^{\prime}m$ qubits. The initial state $|\Psi\rangle$ required for the Grover iteration is obtained by several $H$ transforms. The iterative function $G=D_{|\Psi\rangle}O_{\text{test}}$ is run for $\mathcal{O}(2^{(\kappa-t)/2})$ iterations, where $D_{|\Psi\rangle}=(2|\Psi\rangle\langle\Psi|-I)$. The final step is measuring the index $i^{l}$ to obtain $i_{0}^{l}$. The implementation of $U_{P^{\mathcal{L}}}$ and $O_{\text{test}}$ in Algorithm 4 is similar to that in Ref. [shi2024dedicated]. Specifically, $U_{P^{\mathcal{L}}}$ implements the function
$P^{\mathcal{L}}(i^{l},x^{l})=\bigoplus_{u\in\mathcal{L}}p((x^{l}\oplus i^{l})\|0^{t}\oplus u),$
which is the XOR sum of $2^{t}$ public functions $p$ and can be implemented in parallel without a sequential relationship, where $t>0$ is a tunable truncation parameter.
@@ III-B2 Offline Dedicated Quantum Attacks on Block Ciphers Based on TPP-PRFs
To achieve a usable success probability after $r$ iterations, we need $c^{\prime}=n+\kappa-2t+\pi+1$ queries per iteration. We now give the complexity and success probability of the offline dedicated quantum attack on constructions based on $p$-XOR-type functions in the Q1 (Q2) model in Theorem 4.
Theorem 4
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ be a decoupled XOR-type function, then there exists an offline dedicated quantum attack on constructions based on TPP-PRFs in Algorithm 4. This attack can recover $i_{0}^{l}\in\{0,1\}^{\kappa-t}$ with a time complexity of $O(2^{(\kappa-t)/2}\cdot n^{3})$, requiring $O(2^{n})$ classical queries to the encryption oracle in the Q1 model, and $O(2^{t}\cdot(n-t))$ quantum queries in the Q2 model. The probability of falling below a success rate of
is lower than $2^{n+\kappa-2t-\frac{2^{n-t}}{4c^{\prime}}}$, where $0<t<n$, $c^{\prime}=n+\kappa-2t+\tau+1$.
For $p$-XOR-type functions and their new periodic functions $F^{\mathcal{L}}$ constructed from instantiations of TPP-PRFs in Section III-B1, we can implement offline dedicated attacks to recover the period of $F^{\mathcal{L}}$ by Theorem 4.
@@
7:Apply unitary operation, which writes $\bigoplus\limits_{u\in\mathcal{L}}g_{1}(x^{l}||0^{t}\oplus u)$ in the second register if the first contains the value $x^{l}$:
@@
In the Q1 model, this attack recovers the key $l_{23}^{-1}l_{24}k_{2}^{l}\|l_{13}^{-1}l_{14}k_{1}^{l_{2}}$ of block ciphers based on TPP-PRFs, requiring $O(2^{p+t})$ classical queries to the encryption oracle and $O(2^{(2n-2t-p)/2}\cdot n^{3})$ offline quantum computation time, where 0<p<nt0<p<n-t, $0<t<n$, $l_{2}\in\{0,1\}^{n-t-p}$ and $l\in\{0,1\}^{n-t}$. Let $D$ be the number of classical queries and $T$ be the time of offline quantum computations. This attack achieves a tradeoff of $T^{2}D=2^{2n-t}$, which is balanced at $T=D=\tilde{O}(2^{(2n-t)/3})$. A larger truncation parameter $t$ reduces both the time and the classical query complexity but results in a shorter recovered key length.
@@
The remaining keys can be recovered as follows. We can recover the period $s^{l_{1}}=l_{13}^{-1}l_{14}k_{1}^{l_{1}}$ by solving the equation system $s^{l_{1}}\cdot v_{i}^{l_{1}}=0$, requiring $O(n^{3})$ offline computation time. Subsequently, the remaining $t$-bit can be recovered by applying the offline Simon’s algorithm on $f^{\prime}(k^{r},x)=f(k_{2}^{l}\|i^{r},x)$, which has the period $s$ when $i^{r}=k_{0}^{r}$.
@@
In the Q2 model, our offline attack proposed in Theorem 4 involves the construction of $F^{\mathcal{L}}$ that differs from in Section III-B1 by making it unnecessary to partition the (nt)(n-t)-bit subkey $k_{1}^{l}$ into $k_{1}^{l_{1}}$ with $p$-bit and $k_{1}^{l_{2}}$ with (ntp)(n-t-p)-bit.
TABLE I: Previous and new quantum attacks on $\mathsf{SoEM22}$
Algorithm Model Iteration Query Time Reference
Grover-meets-Simon [leander2017grover] Q2 $2^{n/2}$ $O(2^{n/2}\cdot n)$ $O(2^{n/2}\cdot n^{3})$ [shinagawa2022quantum]
Dedicated attack by constructing an XOR-type function [shi2024dedicated] Q2 $2^{(n-t)/2}$ $O(2^{(n+t)/2}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot n^{3})$ Ours
Offline dedicated attack by constructing a $p$-XOR-type function Q2 $2^{(n-t)/2}$ $O(2^{t}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot n^{3})$ Ours
Offline Simon’s algorithm [bonnetain2019quantum] Q1 $2^{2n/3}$ $O(2^{2n/3})$ $O(2^{2n/3}\cdot n^{3})$ [sun2025quantum]
Offline dedicated attack by constructing a $p$-XOR-type function Q1 $2^{(2n-t)/3}$ $O(2^{(2n-t)/3})$ $O(2^{(2n-t)/3}\cdot n^{3})$ Ours
  • 1
    The fourth column “Query” indicates classical queries to the encryption oracle in the Q1 model and quantum queries in the Q2 model.
  • 2
Taking the $\mathsf{SoEM22}$ construction as an example, we construct a $p$-XOR-type function $f(i,x)=\mathsf{SoEM22}\oplus P_{1}(x)\oplus P_{2}(x\oplus i)$ with the public function $P_{2}(x\oplus i)$, then obtain a new periodic function $F^{\mathcal{L}}(i^{l},x^{l})$ as follows:
$F^{\mathcal{L}}(i^{l},x^{l})=$ $\bigoplus\limits_{u\in\mathcal{L}}\mathsf{SoEM22}(x^{l}\|0^{t}\oplus u)\oplus P_{1}(x^{l}\|0^{t}\oplus u)$ (17)
$\oplus P_{2}\left((x^{l}\oplus i^{l})\|0^{t}\oplus u\right).$
@@
This function has an (nt)(n-t)-bit period $s^{l}=k^{l}_{1}$ when $i^{l}=k^{l}_{2}$. We first prepare the state $|\psi_{G_{1}^{\mathcal{L}}}\rangle$ by requiring $O(2^{t}\cdot(n-t))$ quantum queries to the $\mathsf{SoEM22}$ oracle by Algorithm 7. Then, we apply the offline attack to successfully recover the period $k^{l}_{1}$ with $O(2^{(n-t)/2}\cdot n^{3})$ offline quantum computation time by Algorithm 4. Compared to the attack by Shi et al. [shi2024dedicated], our method reduces the quantum query complexity from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ with the same time complexity, demonstrating the advantage of our offline attack in the Q2 model (See Table I for a comparison with previous attacks and ours on the $\mathsf{SoEM22}$ construction).
IV Conclusion